OpenStudy (anonymous):

Hi..I have a math problem I need help with please... -7Y+30-3(5Y-1)+2Y=8-3Y-(2Y-4)

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

-7y+30-15y+3+2y=8-3y-2y+4 -20y+33=12-5y -15y=-21 y=7/5

OpenStudy (anonymous):

1.2